OpenStudy (anonymous):

can somebody give me the answer to: Solve x + 3y = 9 3x - 3y = -13 Please?

OpenStudy (anonymous):

Add the two equations eliminating Y, and get x.. Then substitute in any of them for the value of x and get the value of y!

OpenStudy (anonymous):

I get to: x = 11/2 But that can't be the right answer :/ I don't know what I'm doing wrong.

OpenStudy (anonymous):

x + 3y = 9 3x - 3y = -13 ------------------ 4x = -4 x = -1

OpenStudy (anonymous):

Do yo see when we add the two equations, the y's cancel. This leaves you with 4x = -4 Divide both sides by 4 to get x = -1

OpenStudy (anonymous):

Now substitute -1 for x in either original equation and solve for y.